OpenStudy (anonymous):

Solve: −3(x − 2) = 15 3 −3 −5.7 −7

OpenStudy (anonymous):

−3(x − 2) = 15 divide by -3 then add two.....

OpenStudy (anonymous):

-3(x-2) = 15 -3x-(-6) = 15 -3x+6 = 15 Now put the 6 to the right by subtracting it from both sides of the equation: -3x = 15 - 6 -3x = 9 Now divide both sides by -3 to find out the value of 'x': x = 9/-3 x = -3 Your answer is: x = -3;)